# Bayesian Optimization in Language Space: An Eval-Efficient AI Self-Improvement Framework

### v2

@@ Appendix C Extended experiment results
Figure 17: T-BoN BO and baseline’s generated Ad images for five fictional brands: Odyssey electric family SUV (Scenario 3), Oasis eco-lodge (Scenario 4), Momentum Digital Banking (Scenario 5), Aeterno luxury watch (Scenario 7), and SyncFlow B2B software (Scenario 8)
Appendix D Proofs
D.1 Auxiliary lemmas
Lemma 3 (Uniform first-order expansion).
Let $g=\nabla\mu(x)$ and $h=\nabla\sigma(x)$. There exist $C>0$ and $\varepsilon_{0}>0$ such that for all $\varepsilon\in(0,\varepsilon_{0}]$ and all $u\in\mathbb{R}^{d}$,
$\mu(x+\varepsilon u)$ $=\mu(x)+\varepsilon\,g^{\top}u\;+\;r_{\mu}(\varepsilon,u),$ $|r_{\mu}(\varepsilon,u)|$ $\leq C\varepsilon^{2},$ (28)
$\sigma(x+\varepsilon u)$ $=\sigma(x)+\varepsilon\,h^{\top}u\;+\;r_{\sigma}(\varepsilon,u),$ $|r_{\sigma}(\varepsilon,u)|$ $\leq C\varepsilon^{2}.$ (29)
Proof.
By the mean-value theorem with Lipschitz gradients, for some $\xi$ between $x$ and $x+\varepsilon u$, $\mu(x+\varepsilon u)=\mu(x)+\nabla\mu(\xi)^{\top}(\varepsilon u).$ Hence $\mu(x+\varepsilon u)=\mu(x)+\varepsilon g^{\top}u+\varepsilon(\nabla\mu(\xi)-\nabla\mu(x))^{\top}u,$ and $|\nabla\mu(\xi)-\nabla\mu(x)|\leq L_{\mu}\|\xi-x\|\leq L_{\mu}\varepsilon$, giving $|r_{\mu}|\leq L_{\mu}\varepsilon^{2}$. The proof for $\sigma$ is identical with $L_{\sigma}$. Taking $C:=\max\{L_{\mu},L_{\sigma}\}$ yields the stated bounds with the same $C$. ∎
Lemma 4 (Max-stability under bounded perturbations).
$g_{i_{f}}\geq f_{i_{f}}-\Delta\quad\text{and}\quad f_{i_{g}}\geq g_{i_{g}}-\Delta,$ (30)
Proof.
Lemma 5 (Spherical cap coverage).
$\mathbb{P}\!\left(\max_{i\leq N}v^{\top}U_{i}\geq 1-\eta_{N}\right)\to 1.$ (31)
Proof.
Lemma 6 (Near-maximum coupling in a thin band).
Let $M_{N}=\max_{i}\xi_{i}$ and define $v_{N}:=(g+M_{N}h)/\|g+M_{N}h\|$. Also, $S_{i}:=a_{i}+b_{i}M_{N}$, $Y_{i}^{(0)}:=a_{i}+b_{i}\xi_{i}$, $b_{i}:=\sigma(x)+\varepsilon h^{\top}U_{i}$. Fix any sequence $c_{N}\uparrow\infty$ with $c_{N}/q_{N}\to 0$ and set $\delta_{N}:=c_{N}/q_{N}$. Then with probability $\to 1$ there exists an index $j$ such that
$\xi_{j}\geq M_{N}-\delta_{N}\quad\text{and}\quad v_{N}^{\top}U_{j}\geq 1-\eta_{N},$ (32)
whenever $\eta_{N}\downarrow 0$ satisfies
$N\,\mathbb{P}(\xi\geq M_{N}-\delta_{N})\,\rho\{u:v_{N}^{\top}u\geq 1-\eta_{N}\}\to\infty.$ (33)
Moreover, for this $j$,
$S_{j}-Y_{j}^{(0)}=b_{j}(M_{N}-\xi_{j})\leq b_{\max}\,\delta_{N}.$ (34)
Proof.
$E_{N}:=\{M_{N}\geq q_{N}-\delta_{N}\}.$ (35)
$\mathbb{P}(\xi\geq t_{N})\;\geq\;\mathbb{P}(\xi\geq M_{N}-\delta_{N})\quad\text{on }E_{N}.$ (36)
Let
$K_{N}:=\sum_{i=1}^{N}\mathbf{1}\{\xi_{i}\geq t_{N}\}\sim\mathrm{Binomial}\!\big(N,\,p_{N}\big),\qquad p_{N}:=\mathbb{P}(\xi\geq t_{N}).$ (37)
A Chernoff bound gives
$\mathbb{P}\!\left(K_{N}<\tfrac{1}{2}Np_{N}\right)\;\leq\;\exp\!\left(-\tfrac{1}{8}Np_{N}\right).$ (38)
Consider
$L_{N}\;:=\;\sum_{i=1}^{N}\mathbf{1}\{\xi_{i}\geq t_{N}\}\,\mathbf{1}\{v_{N}^{\top}U_{i}\geq 1-\eta_{N}\}.$ (39)
[…]
$L_{N}\,\big|\,\{\xi_{i}\}\;\sim\;\mathrm{Binomial}\!\big(K_{N},\,p_{\mathrm{cap},N}\big),\qquad\mathbb{P}\big(L_{N}=0\,\big|\,\{\xi_{i}\}\big)\;=\;(1-p_{\mathrm{cap},N})^{K_{N}}\;\leq\;\exp\!\big(-K_{N}p_{\mathrm{cap},N}\big).$ (40)
$\leq\mathbb{E}\!\left[\exp\!\big(-\tfrac{1}{2}Np_{N}\,p_{\mathrm{cap},N}\big)\right]\;+\;\exp\!\big(-\tfrac{1}{8}Np_{N}\big)\;+\;o(1).$ (41)
On $E_{N}$, (36) yields $p_{N}\geq\mathbb{P}(\xi\geq M_{N}-\delta_{N})$. By the lemma’s hypothesis,
Since $0\leq e^{-\frac{1}{2}Z_{N}}\leq 1$, for any fixed $T>0$,
Letting $T\to\infty$ gives $\mathbb{E}[e^{-\frac{1}{2}Z_{N}}]\to 0$. Moreover, $Z_{N}\leq Np_{N}$ since $p_{\mathrm{cap},N}\leq 1$. If $Np_{N}$ were bounded along a subsequence, then $Z_{N}$ would be bounded along that subsequence, contradicting $Z_{N}\xrightarrow{p}\infty$. Hence $Np_{N}\to\infty$, so $\exp(-\tfrac{1}{8}Np_{N})\to 0$. Therefore the second term in (41) vanishes, and $\mathbb{P}(L_{N}=0)\to 0$.
D.2 Proof of Theorem 2
Proof.
By Lemma 3,
$Y_{i}$ $=\mu(x)+\varepsilon g^{\top}u_{i}+r_{\mu}(\varepsilon,u_{i})+\big(\sigma(x)+\varepsilon h^{\top}u_{i}+r_{\sigma}(\varepsilon,u_{i})\big)\xi_{i}$
$=a_{i}+b_{i}\xi_{i}+R_{i},$ (42)
where $a_{i}:=\mu(x)+\varepsilon g^{\top}u_{i}$, $b_{i}:=\sigma(x)+\varepsilon h^{\top}u_{i}$, and $R_{i}:=r_{\mu}(\varepsilon,u_{i})+\xi_{i}r_{\sigma}(\varepsilon,u_{i}).$ Then $|R_{i}|\leq C\varepsilon^{2}(1+|\xi_{i}|)$ for all $i$. Define $S_{i}:=a_{i}+b_{i}M_{N}$, $b_{\max}:=\max_{1\leq i\leq N}|b_{i}|$, and $U_{i}:=u_{i}$.
Now let $Y_{i}^{(0)}:=a_{i}+b_{i}\xi_{i}$ and $i_{0}\in\arg\max_{i}Y_{i}^{(0)}$. From (42),
By the bound on $R_{i}$ and $|\xi_{i}|\leq M_{N}:=\max_{j}\xi_{j}$,
$Y_{i^{\star}}^{(0)}\geq\max_{i}Y_{i}^{(0)}-2C\varepsilon^{2}(1+M_{N}).$ (43)
Also,
where the last term is $\leq 0$ provided $b_{i}\geq 0$ for all $i$, which holds by Assumption A4 for $\varepsilon\leq\varepsilon_{0}:=\sigma(x)/(2\|h\|)$ when $\sigma(x)>0$ since $b_{i}=\sigma(x)+\varepsilon h^{\top}U_{i}\geq\sigma(x)-\varepsilon\|h\|\geq\tfrac{1}{2}\sigma(x)>0$. Therefore,
$\max_{i\leq N}S_{i}-S_{i^{\star}}\;\leq\;\varepsilon\|g+M_{N}h\|\,\eta_{N}\;+\;b_{\max}\delta_{N}\;+\;4C\varepsilon^{2}(1+M_{N}).$ (44)
Set:
By Lemma 3,
$S_{i}\;=\;\mu(x)+M_{N}\sigma(x)+\varepsilon(g+M_{N}h)^{\top}u_{i}+\tilde{r}_{i},\qquad|\tilde{r}_{i}|\leq C\varepsilon^{2}(1+M_{N}).$
Hence for any $i$,
so
$\big|\,(S_{i}-S_{i^{\star}})-\varepsilon\|g+M_{N}h\|\,(v_{N}^{\top}U_{i}-v_{N}^{\top}\widehat{u}_{N})\,\big|\;\leq\;2C\varepsilon^{2}(1+M_{N}).$ (45)
Taking $i$ that attains $\max_{i}S_{i}$ and using (44) yields
$\max_{i\leq N}v_{N}^{\top}U_{i}-v_{N}^{\top}\widehat{u}_{N}\;\leq\;\eta_{N}\;+\;\frac{\Delta^{\prime}_{N}(\varepsilon)}{\varepsilon\|g+M_{N}h\|}.$ (46)
Thus
[…]
Assume $\|h\|>0$ (the $\|h\|=0$ case is addressed in Remark 1). Using
$\|g+M_{N}h\|\;\geq\;M_{N}\|h\|-\|g\|$
and Lemma 1, $M_{N}\to\infty$ and $M_{N}\asymp q_{N}$. Also $b_{\max}\leq\sigma(x)+\varepsilon\|h\|+C\varepsilon^{2}$. Hence
$N\,e^{c_{N}}\,\rho\{u:v_{N}^{\top}u\geq 1-\eta_{N}\}\to\infty$
makes the RHS of (47) equal to $o_{p}(1)+O(\varepsilon)$ for fixed $\varepsilon>0$; then let $\varepsilon\downarrow 0$.
With $N\to\infty$ first and fixed $\varepsilon>0$, (47) implies $1-v_{N}^{\top}\widehat{u}_{N}\xrightarrow{p}0$. Then $\varepsilon\downarrow 0$ removes the $O(\varepsilon)$ residual, yielding $v_{N}^{\top}\widehat{u}_{N}\xrightarrow{p}1$. Since $M_{N}-q_{N}\to 0$ in probability by Lemma 1, the unit vectors $v_{N}=(g+M_{N}h)/\|g+M_{N}h\|$ and $\tilde{v}_{N}:=(g+q_{N}h)/\|g+q_{N}h\|$ satisfy $\|v_{N}-\tilde{v}_{N}\|\to 0$ in probability. Hence $\widehat{u}_{N}\xrightarrow{p}\tilde{v}_{N}$, which is the claimed gradient direction of $A_{\beta_{N}}$ with $\beta_{N}=q_{N}$.
By Lemma 3,
$A_{\beta_{N}}(x+\Delta x^{(N)})=A_{\beta_{N}}(x)+\varepsilon\nabla A_{\beta_{N}}(x)^{\top}\widehat{u}_{N}+O(\varepsilon^{2}).$
Write $\nabla A_{\beta_{N}}(x)=g+q_{N}h$. Using $\tilde{v}_{N}^{\top}\widehat{u}_{N}\xrightarrow{p}1$ gives $\nabla A_{\beta_{N}}(x)^{\top}\widehat{u}_{N}=\|\nabla A_{\beta_{N}}(x)\|\,\tilde{v}_{N}^{\top}\widehat{u}_{N}=\|\nabla A_{\beta_{N}}(x)\|(1-o_{p}(1)).$ Thus $A_{\beta_{N}}(x+\Delta x^{(N)})\geq A_{\beta_{N}}(x)+\varepsilon\|\nabla A_{\beta_{N}}(x)\|-o_{p}(\varepsilon),$ as claimed. ∎
Remark 1 (Edge cases).
(i) If $\|h\|=0$, then $\nabla A_{\beta_{N}}(x)=g$ and the direction reduces to $\nabla\mu(x)$. (ii) If $\sigma(x)=0$ but $\|h\|>0$, the result holds unchanged; only the linearization constants change. (iii) If $\sigma(x)=0$ and $\|h\|=0$, the model is locally deterministic and the problem is degenerate; selection aligns with $\nabla\mu(x)$ or is undefined if $g=0$.
Appendix E GP-UCB and its evaluation efficiency
Setup.
[…]
The objectives.
$\mathcal{R}_{C}(T;f,\mathcal{A}):=\sum_{t=1}^{T}\left[\max_{\mathbf{x}\in\mathcal{D}}f(\mathbf{x})-f\left(\mathbf{x}_{t}\right)\right]$ (48)
The other is simple regret, defined as
$\mathcal{R}_{S}(T;f,\mathcal{A}):=\max_{\mathbf{x}\in\mathcal{D}}f(\mathbf{x})-f\left(\mathbf{x}^{(T)}\right)$ (49)
The lower bound.
$\inf_{\mathcal{A}}\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}\mathbb{E}\left[\mathcal{R}_{C}(T;f,\mathcal{A})\right]=\begin{cases}\Omega(T^{\frac{\nu+d}{2\nu+d}}),&\text{ for Matérn kernels, }\\ \Omega(T^{\frac{1}{2}}\ln^{\frac{d}{2}}(T)),&\text{ for SE kernels. }\end{cases}$ (50)
$\inf_{\mathcal{A}}\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}T\left(\epsilon,f,\mathcal{A}\right)=\begin{cases}\Omega\left(\left(\frac{1}{\epsilon}\right)^{2+d/\nu}\right),&\text{ for Matérn kernels, }\\ \Omega\left(\frac{1}{\epsilon^{2}}\left(\log\frac{1}{\epsilon}\right)^{d/2}\right),&\text{ for SE kernels. }\end{cases}$ (51)
where Matérn kernels are defined as
$\Psi_{\mathrm{M}}\left(\mathbf{x}-\mathbf{x}^{\prime}\right):=\frac{1}{\Gamma(\nu)2^{\nu-1}}\left(\frac{2\sqrt{\nu}\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}}{\ell}\right)^{\nu}K_{\nu}\left(\frac{2\sqrt{\nu}\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}}{\ell}\right),\quad\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{D}$ (52)
and SE kernels are defined as
$\Psi_{\mathrm{SE}}\left(\mathbf{x}-\mathbf{x}^{\prime}\right)=\exp\left(-\frac{\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}^{2}}{2\ell^{2}}\right),\quad\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{D}$ (53)
GP-UCB algorithm.
$\mathbb{E}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]=\mathbf{k}_{t}^{\top}(\mathbf{x})\left(\mathbf{K}_{t}+\sigma^{2}\mathbf{I}_{t}\right)^{-1}\mathbf{y}_{t},\operatorname{Var}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]=k(\mathbf{x},\mathbf{x})-\mathbf{k}_{t}(\mathbf{x})^{\top}\left(\mathbf{K}_{t}+\sigma^{2}\mathbf{I}_{t}\right)^{-1}\mathbf{k}_{t}(\mathbf{x}),$ (54)
The evaluation efficiency of GP-UCB.
whitehouse2023sublinear, wang2023regret showed that GP-UCB srinivas2012information upper bound on the cumulative regret of any algorithm for $f\in\mathcal{N}_{\Psi}(\mathcal{D})$ that is almost nearly tight. Specifically, for any constant $B>0$, the GP-UCB algorithm $\mathcal{A}_{\text{GPUCB}}$ achieves
$\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}\mathbb{E}\left[\mathcal{R}_{C}(T;f,\mathcal{A}_{\text{GPUCB}})\right]=\begin{cases}O(T^{\frac{\nu+d}{2\nu+d}}),&\text{ for Matérn kernels, }\\ O(T^{\frac{1}{2}}\ln^{\frac{d}{2}}(T))&\text{ for SE kernels. }\end{cases}$ (55)
$\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}T(\epsilon,f,\mathcal{A}_{\mathrm{GPUCB}})=\begin{cases}O\left(\left(\frac{1}{\epsilon}\right)^{2+d/\nu}\right),&\text{for Matérn kernels,}\\ O\left(\frac{1}{\epsilon^{2}}\cdot\left(\log\frac{1}{\epsilon}\right)^{d+3}\right),&\text{for SE kernels.}\end{cases}$ (56)

### v3

@@ D.2 Detailed Experiment Results
Figure A8: TextBO and baseline’s generated Ad images for five fictional brands: Odyssey electric family SUV (Scenario 3), Oasis eco-lodge (Scenario 4), Momentum Digital Banking (Scenario 5), Aeterno luxury watch (Scenario 7), and SyncFlow B2B software (Scenario 8)
D.3 Final Prompts in TextBO and GEPA
Prompt for generating the TextBO $(T=10)$ ad for GreenBite (plant-based burger patty) "**Key Message:**
*Core Components:**
**Stylistic Qualities:**
Figure A9: Prompts for TextBO ($T=10$)
**Core Components:**
Figure A10: Prompts for GEPA ($T=10$)
Appendix E Appendix for Section 9
E.1 Implementation Details of GEPA and TextBO-GEPA for agentic AI experiments
GEPA
GEPA [agrawal2025gepa] is an iterative, prompt-optimization-based, self-improving AI method. The key idea of GEPA is to “combine reflection with evolution”. Here, an evolution is the same concept as taking a textual gradient in textgrad. GEPA iterates over three steps in each optimization iteration: Step 1) choosing a prompt and evaluating it; Step 2) updating the meta-reflection; and Step 3) improving the prompt.
For Step 1), it uses instance-level information to (i) maintain a Pareto archive of which prompts in the candidate prompt set currently win on which problem instances, (ii) stochastically sample a non-dominated candidate prompt (favoring candidates that win on more instances), and (iii) evaluate that sampled candidate by rolling it out on a minibatch of problem instances.
In Step 2), GEPA updates a global meta-reflection $R_{t}$ that summarizes what has tended to work well or poorly so far. Concretely, after collecting rollouts and their scores, it applies a reflection operator to the accumulated history, $R_{t}:=\texttt{MetaReflect}(H_{t};M_{\text{critic}})$, where the critic is shown a context-limited subset of representative high-performing and low-performing rollouts (including the prompts and outcomes) and asked to (i) identify recurring success patterns and failure modes and (ii) distill them into a short list of actionable, prompt-level guidelines. This shared reflection is carried forward and provided as context to subsequent prompt-evolution (textual-gradient) steps. This step is fundamentally the same as TextBO’s meta-reflection steps.
For Step 3), we take one step of evolution (or equivalently, one step of textual gradient) of the prompt we evaluated in Step 1) and accept it only if it has a better evaluation than the original prompt. When the offspring prompt is accepted, we append it to the candidate prompt set without removing the original prompt from the candidate set.
TextBO-GEPA
TextBO-GEPA is identical to GEPA except for the prompt-improvement step (Step 3): whenever GEPA would take a single evolution step (one textual-gradient update) to propose an offspring prompt, TextBO-GEPA instead performs multiple successive Best-of-N textual-gradient steps (as in §5) before evaluating and applying the same acceptance rule as GEPA. That is, at each inner step, it samples $N$ local textual edits, applies them to form $N$ candidate prompts, generates the corresponding candidate outcomes, and uses the critic (conditioned on the current reflection) to select the most promising candidate; repeating this for several steps yields a refined offspring prompt, which is then accepted and appended to the candidate prompt set only when it improves upon its parent under the evaluation procedure.